A 3250 kg rocket is traveling through space at 6.70 m/s. Thrusters are fired for 3.10 seconds. This causes the rocket to increase its velocity to 10.5 m/s. Calculate the average force needed to cause this increase.
1
Expert's answer
2020-11-02T09:24:48-0500

We can find the average force needed to cause the increase of rocket's velocity from the

impulse-momentum change equation:


"F_{avg}\Delta t=m\Delta v,""F_{avg}\Delta t=m(v_f-v_i),"

here, "F_{avg}" is the average force needed to cause the increase of rocket's velocity,

"\Delta t=3.10\ s" is time during which the thrusters are fired, "m=3250\ kg" is the mass of the rocket, "v_i=6.70\ \dfrac{m}{s}" is the initial velocity of the rocket, "v_f=10.5\ \dfrac{m}{s}" is the final velocity

of the rocket.

Then, from this equation we can calculate "F_{avg}":


"F_{avg}=\dfrac{m(v_f-v_i)}{\Delta t},""F_{avg}=\dfrac{3250\ kg\cdot(10.5\ \dfrac{m}{s}-6.70\ \dfrac{m}{s})}{3.10\ s}=3984\ N."

Answer:

"F_{avg}=3984\ N."